-20 .22 10) - 72 +16 A) {0} 10) B) (-60) C) (-0,0,0,0(0,0) D) (-0,4) (4)
Answer #1

10)

\small \frac{x^{2}}{x^{2}+16}\geq 0

Numerator and Denominator is always positive for any real number.

Therefore, the function is always positive for any real number.

\small \boldsymbol{\therefore x\in (-\infty ,\infty )}